Shift report 20000813 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 13 2000 - 15:57:03 EDT

  • Next message: http account: "Shift report 20000814 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Aug 13 15:57:03 2000, Shift: 08:00-16:00
    Supervisor: FR, with: TF
    Log book#4 pages 79 - 81
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 14:00. we got a good store with coggeg beam at 
     gamma=70. It is still going. Beam intensities: b=6x10**9 
     and y=3x10**9.
    
    2. Run 2235: MRS at 90deg. with D5 off, about 24423 coll.
     Run 2236: MRS at 90deg. with D5 on (700A, 4kGauss), 
     still running.
    
    3. T2 readout board still broken. 
     No other detector problems.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 13 2000 - 16:01:21 EDT